Two point charges lie on the x-axis: 1.70 μC at x = 1.00 m and –2.20 μC at x = 2.35 m .

Find the place(s) where the electric field is zero.
Answer #1

Solution)

We know,

1/4*pi*Eo *(q1/r1^2) =1/4*pi*Eo *(q2/r2^2)

1.7/(x-1)^2 =2.2/(2.35-x)^2

On solving, we get

X=2.36 m (Ans)

======